# I2CKD : Intra- and Inter-Class Knowledge Distillation for Semantic Segmentation

### v1

@@ Abstract
Index Terms—  Knowledge distillation, semantic segmentation, deep learning, class prototype, triplet loss
@@ 1 Introduction
In recent years, deep neural networks have achieved impressive performance in semantic segmentation. However, this good performance comes at the expense of heavy storage and computation, which impedes their deployment on resource-limited edge devices. To address this issue, some compact segmentation networks are proposed such as ESPNet [2], ENet [3] and BiseNet [4]. Another category of approaches tries to compress the existing cumbersome models, which can be categorized into pruning [5], quantization [6], and knowledge distillation (KD) [7]. Pruning-based methods strive to reduce the network size by eliminating unnecessary connections between neurons in adjacent layers. Quantization seeks to decrease the number of bits per network weight by substituting data types, such as using 8-bit integers instead of 32-bit floating-point numbers. In both the pruning and quantization schemes, the compressed network is trained independently of the cumbersome one which affect drastically the good performance. As a promising mechanism, KD [7] transfer the knowledge from a cumbersome network to a compact one. The former is called teacher and the latter is called student. Since the KD trains a student with the guidance of the teacher, it achieves impressive results in classification [8] and consequently is extended to the semantic segmentation task. In this context, the common point of the proposed methods in the literature is the exploitation of the knowledge at feature maps levels because semantic segmentation needs a structured output. In order to accomplish this, the state-of-the-art focus on developing new pixel-wise [9, 10, 11, 12], channel-wise [13, 14, 15] or image-wise [16] correlations between teacher and student feature maps. In addition, the ground-truth and the knowledge transferred at score maps level is often taken into consideration. Generally, the quality of the knowledge between the teacher and student is computed in densely pairwise manner using several metrics such as Mean Square Error (MSE) without consideration of the intra- and inter-class features relation.
@@
Fig. 1: Our motivation is to minimize the intra-class distances between teacher and student prototypes of the same class ($p_{c}^{t}$ and $p_{c}^{s}$) and to maximize the inter-class distances between teacher and student prototypes of the different classes ($p_{j}^{t}$ $\forall j\neq c$ and $p_{c}^{s}$). The dashed arrow refers to distance.
@@ 2.2 Knowledge distillation for semantic segmentation
The Fully Convolutional Networks [23] such as PSPNet [24] and DeepLab [25] have dominated the leaderboard on several semantic segmentation dataset. Inspiring by the success of KD in image classification, some methods rely on this technique to train compact FCN networks by capturing the correlation information among pixels, channels and images. In [9], Xie et al. extracts two types of knowledge: zero order and first order. The zero order computes the difference between the pixels class probabilities. Regarding to the first order, the difference between the center pixel and its $8$-neighborhood is considered. He et al. [12] focus on matching the feature maps size of teacher and student, and propose to apply an auto-encoder. After that, a pairwise affinity map is computed to quantify the relationship between the teacher and student knowledge. Liu et al. [10] propose a Structured Knowledge Distillation (SKD) scheme by considering the intermediate distillation between feature maps using a graph, the pixel distillation between score maps, and holistic distillation through adversarial learning. Wang et al. [11] develop a new method called Intra-class Feature Variation Distillation (IFVD) that learn the student to mimic the intra-class relation of the teacher. To transfer the IFV knowledge, the cosine distance is used. Shu et al. [14] develop a method called Channel-Wise knowledge Distillation (CWD) that compute the KLD between the softmax of the teacher and student activation channel maps along the channel dimension. In our previous work [26], we leverage the interdepencies between teacher feature maps as knowledge to distill. This knowledge is captured using the self-attention mechanism and transferred using the MSE.
@@ 3.1 Overview
As illustrated in the Figure 2, our method I2CKD distills the knowledge from a well-trained teacher network to a student. For each training epoch of student network, we froze the teacher network. The student network updates its weights through three losses that capture, respectively, the difference student/ground-truth and the difference teacher/student at score maps and feature maps levels. The core of the contribution lies at the feature maps level where we propose to exploit the intra- and inter-class relations between teacher and student. For doing so, we compute the triplet loss between the teacher and the student class prototypes.
Fig. 2: Flowchart of the proposed method: I2CKD. The student network is trained via three losses: (1) a cross-entropy loss $\mathcal{L}_{task}$, (2) a KLD between score maps $\mathcal{L}_{SM}$ , and (3) a triplet loss between prototype classes $\mathcal{L}_{I2CKD}$. At feature maps position, we aim to learn the student network to mimic the intra- (to minimize) and the inter-class (to maximize) relations of the teacher through a triplet loss. We underline that the teacher network is frozen during the student network training.
@@ 3.2.2 Triplet loss
The enforcement of this constraint allows to formulate the loss between student and teacher class prototypes as follows:
@@ 3.3 Overall student training loss
    where $\phi(\cdot)$ is the softmax function:
@@ 4.2 Ablation study
We recall that the student network is trained using the general loss expressed in Equation 4 which contains three losses. In Table 1, we conduct an experiment to study their contribution on the validation set of Cityscapes. Based on the results presented in this table, the following observations can be made: 1) the distillation at score maps brings more gains than the original student. Further aligning the proposed class prototypes distillation using triplet loss boosts the performance and brings the most performance gains.
@@
$\mathcal{L}_{seg}$ $\mathcal{L}_{SM}$ $\mathcal{L}_{FM}$
$\checkmark$ 73.21
@@ 4.3.1 Comparison on Cityscapes
In Table 2, we compare the proposed method with contemporary distillation methods tailored to semantic segmentation including SKD [10], IFVD [11], and CWD [14]. To ensure a fair comparison, we adopt different teacher-student network pairs by changing the segmentation framework or backbones. An intuitive idea for compressing a heavy FCN is to replace its backbone by a compact one. This table demonstrates that replacing the ResNet101 backbone of DL model with ResNet18 resulted in a decrease in segmentation performance. Specifically, in the validation set, the performance dropped from $78.07\%$ to $74.21\%$. The decline in performance is primarily due to the reduced size of the student model ($13.6$M) which make an independent training without direct guidance from the teacher model. Additionally, it is noteworthy that both the proposed method and all the compared distillation methods greatly improve the performance of the student networks which demonstrate the interest of adopting knowledge distillation for neural network compression. From this table, we can see that our method I2CKD significantly outperforms the compared distillation methods when we consider a student with the same segmentation framework of the teacher (DL-R101/DL-R18) as well as when we use a student with a different segmentation framework (DL-R101/PSPNet-R18). The aforementioned results strongly demonstrate the benefit of using the class prototypes for knowledge extraction and triplet loss for knowledge transfer. Some qualitative results using the configuration DL-R101/DL-R18 are illustrated in Figure 3.
Teacher Methods Params val $mIoU$
@@
Teacher DL-R101 61.1 78.07
DL-R18 13.6 74.21
DL-R101 SKD [10] 75.42
IFVD [11] 75.59
DL-R18 CWD [14] 13.6 75.55
I2CKD (ours) 76.07
PSPNet-R18 12.9 72.55
DL-R101 SKD [10] 73.29
IFVD [11] 73.71
PSPNet-R18 CWD [14] 12.9 74.36
I2CKD (ours) 74.43
Table 2: Quantitative results on Cityscapes.
Fig. 3: Qualitative segmentation results on Cityscapes validation set using DeepLabV3-R18 student.
@@ 4.3.2 Comparison on Pascal VOC
We additionally perform experiments on the PASCAL VOC dataset to further validate the distillation capability of the proposed I2CKD in visual object segmentation. The results are recorded in Table 3. Both original student models drop the performance of the teacher one (cumbersome) by $4.46\%$ and $4.43\%$ respectively. While the other distillation methods only bring limited improvements, our I2CKD achieve the best validation $mIoU$ of $76.31\%$ and $74.59\%$ using DL-R18 and PSPNet-R18 students respectively.
@@
Table 3: Quantitative results on Pascal VOC.
@@ 4.3.3 Comparison on CamVid
Table 4: Quantitative results on CamVid.

### v2

@@ Abstract
keywords:
Knowledge distillation, semantic segmentation, deep learning, class prototype, triplet loss
journal: Computer Vision and Image Understanding
@@ 1 Introduction
In recent years, deep neural networks have achieved impressive performance in semantic segmentation. However, this good performance comes at the expense of heavy storage and computation, which impedes their deployment on resource-limited edge devices. To address this issue, some compact segmentation networks are proposed such as ESPNet [2], ENet [3] and BiseNet [4]. Another category of approaches tries to compress the existing cumbersome models, which can be categorized into pruning [5], quantization [6], and knowledge distillation (KD) [7]. Pruning-based methods strive to reduce the network size by eliminating unnecessary connections between neurons in adjacent layers. Quantization seeks to decrease the number of bits per network weight by substituting data types, such as using 8-bit integers instead of 32-bit floating-point numbers. In both the pruning and quantization schemes, the compressed network is trained independently of the cumbersome one which affect drastically the good performance. As a promising mechanism, KD [7] transfers the knowledge from a cumbersome network to a compact one. The former is called teacher and the latter is called student. Since the KD trains a student with the guidance of the teacher, it achieves impressive results in classification [8] and consequently is extended to the semantic segmentation task. In this context, the common point of the proposed methods in the literature is the exploitation of the knowledge at feature maps levels because semantic segmentation needs a structured output. In order to accomplish this, the state-of-the-art focus on developing new pixel-wise [9, 10, 11, 12], channel-wise [13, 14, 15] or image-wise [16] correlations between teacher and student feature maps. In addition, the ground-truth and the knowledge transferred at score maps level are often taken into consideration. Generally, the quality of the knowledge between the teacher and student is computed in densely pairwise manner using several metrics such as Mean Squared Error (MSE) without consideration of the intra- and inter-class features relation.
@@
Figure 1: Our motivation is to minimize the intra-class distances between teacher and student prototypes of the same class ($p_{c}^{t}$ and $p_{c}^{s}$) and to maximize the inter-class distances between teacher and student prototypes of the different classes ($p_{j}^{t}$ $\forall j\neq c$ and $p_{c}^{s}$). The dashed arrow refers to distance.
@@ 2.2 Knowledge distillation for semantic segmentation
The Fully Convolutional Networks [23] such as PSPNet [24] and DeepLab [25] have dominated the leaderboard on several semantic segmentation dataset. Inspiring by the success of KD in image classification, some methods rely on this technique to train compact FCN networks by capturing the correlation information among pixels, channels and images. In [9], Xie et al. extracts two types of knowledge: zero order and first order. The zero order computes the difference between the pixels class probabilities. Regarding to the first order, the difference between the center pixel and its $8$-neighborhood is considered. He et al. [12] focus on matching the feature maps size of teacher and student, and propose to apply an auto-encoder. After that, a pairwise affinity map is computed to quantify the relationship between the teacher and student knowledge. Liu et al. [10] propose a Structured Knowledge Distillation (SKD) scheme by considering the intermediate distillation between feature maps using a graph, the pixel distillation between score maps, and holistic distillation through adversarial learning. Wang et al. [11] develop a new method called Intra-class Feature Variation Distillation (IFVD) that learn the student to mimic the intra-class relation of the teacher. To transfer the IFV knowledge, the cosine distance is used. Shu et al. [14] develop a method called Channel-Wise knowledge Distillation (CWD) that compute the KLD between the softmax of the teacher and student activation channel maps along the channel dimension. Yang et al. [16] proposed new method named CIRKD that capture two types of knowledge: pixel-to-pixel (intra-image) and pixel-to-region (cross-image). Unlike previous methods, Qiu et al. [26] exploit the labels as a source of knowledge.
@@ 3.1 Overview
As illustrated in the Figure 2, our method I2CKD distills the knowledge from a well-trained teacher network to a student. For each training epoch of student network, we froze the teacher network. The student network updates its weights through three losses that capture, respectively, the difference student/ground-truth labels and the difference teacher/student at score maps and feature maps levels. The core of the contribution lies at the feature maps level where we propose to exploit the intra- and inter-class relations between teacher and student. For doing so, we compute the triplet loss between the teacher and the student class prototypes.
Figure 2: Flowchart of the proposed method: I2CKD. The student network is trained via three losses: (1) a cross-entropy loss $\mathcal{L}_{task}$, (2) a KLD between score maps $\mathcal{L}_{SM}$ , and (3) a triplet loss between prototype classes $\mathcal{L}_{I2CKD}$. At feature maps position, we aim to learn the student network to mimic the intra- (to minimize) and the inter-class (to maximize) relations of the teacher through a triplet loss. We underline that the teacher network is frozen during the student network training.
@@ 3.2.2 Triplet loss
The enforcement of the constraint of Equation 2 allows to formulate the loss between student and teacher class prototypes as follows:
@@ 3.3 Overall student training loss
    where $H$ and $W$ denote the height and width respectively. $\phi(\cdot)$ is the softmax function:
@@ 4.1 Experimental setup
In order to make a fair comparison, we follow the same training protocol as the CIRKD method [16] and we use their codebase.
@@ 4.2 Ablation study
We recall that the student network is trained using the general loss expressed in Equation 4 which contains three losses. In Table 1, we conduct an experiment to study their contribution on the validation set of Cityscapes. Based on the results presented in this table, the following observations can be made: 1) The distillation at score maps brings more gains than the original student. 2) Further aligning the proposed class prototypes distillation using triplet loss boosts the performance and brings the most performance gains.
@@
$\mathcal{L}_{seg}$ $\mathcal{L}_{SM}$ $\mathcal{L}_{I2CKD}$
$\checkmark$ 74.21
@@ 4.3.1 Comparison on Cityscapes
In Table 2, we compare the proposed method with contemporary distillation methods tailored to semantic segmentation including SKD [10], IFVD [11], and CWD [14]. To ensure a fair comparison, we adopt different teacher-student network pairs by changing the segmentation framework or backbones. An intuitive idea for compressing a heavy FCN is to replace its backbone by a compact one. This table demonstrates that replacing the ResNet101 backbone of DL model with ResNet18 resulted in a decrease in segmentation performance. Specifically, in the validation set, the performance dropped from $78.07\%$ to $74.21\%$. The decline in performance is primarily due to the reduced size of the student model ($13.6$M) which make an independent training without direct guidance from the teacher model. Additionally, it is noteworthy that both the proposed method and all the compared distillation methods greatly improve the performance of the student networks which demonstrate the interest of adopting knowledge distillation for neural network compression.
From this table, we can see that our method I2CKD significantly outperforms the compared distillation methods [10, 11, 14, 16] and presents competitive results with CIRKD [16] when we consider a pre-trained student on ImageNet with the same segmentation framework of the teacher (DL-R101/DL-R18) as well as when we use a student with a different segmentation framework (DL-R101/PSPNet-R18). We report also in this table the results for a student with ResNet18 backbone trained from scratch (DL-R18*). In this case, the proposed method outperforms the best competitive method CIRKD [16] by $4.93\%$ $mIoU$ on the validation set and $3.53\%$ on test $mIoU$.
The aforementioned results strongly demonstrate the benefit of using the class prototypes for knowledge extraction and triplet loss for knowledge transfer. Some qualitative results of out method on DL-R101/DL-R18 configuration are illustrated in Figure 3.
Teacher Methods Params val $mIoU$ test $mIoU$
@@
Teacher DL-R101 61.1 78.07 77.46
DL-R18 13.6 74.21 73.45
DL-R101 SKD [10] 75.42 74.06
IFVD [11] 75.59 74.26
DL-R18 CWD [14] 13.6 75.55 74.07
CIRKD [16] 76.38 75.05
I2CKD (ours) 76.07 74.66
DL-R18* 13.6 65.17 65.47
DL-R101 SKD [10] 67.08 66.71
IFVD [11] 65.96 65.78
DL-R18* CWD [14] 13.6 67.74 67.35
CIRKD [16] 68.18 68.22
I2CKD (ours) 73.11 71.75
PSPNet-R18 12.9 72.55 72.29
DL-R101 SKD [10] 73.29 72.95
IFVD [11] 73.71 72.83
PSPNet-R18 CWD [14] 12.9 74.36 73.57
CIRKD [16] 74.73 74.05
I2CKD (ours) 74.43 73.97
Table 2: Quantitative segmentation results on Cityscapes. * indicates that the models are trained from scratch. The best result is in bold, and the second-best result is underlined.
Figure 3: Qualitative segmentation results on Cityscapes validation set using DeepLabV3-R18 student.
@@ 4.3.2 Comparison on Pascal VOC
We additionally perform experiments on the PascalVOC dataset to further validate the distillation capability of the proposed I2CKD in visual object segmentation. The results are recorded in Table 3. Both original student models drop the performance of the teacher one (cumbersome) by $4.46\%$ and $4.43\%$ respectively. Our method demonstrates high competitive segmentation results in comparison with the other methods.
@@
Table 3: Quantitative segmentation results on Pascal VOC. The best result is in bold, and the second-best result is underlined.
@@ 4.3.3 Comparison on CamVid
Table 4: Quantitative segmentation results on CamVid. The best result is in bold, and the second-best result is underlined.